OpenStudy (anonymous):

Suppose you have three consecutive integers. The greatest of the three integers is twice as great as the sum of the first two. What are the integers?

OpenStudy (anonymous):

n n+1 n+2 =2 (n+n+1) n+2 = 2(2n +1) n+2 = 4n +2 -3n = 0 n = 0 so the integers are 0,1,2

OpenStudy (anonymous):

let me know if you need more explanation